### v2

@@ IV-B Subtext Inference
We designed multiple metrics to enable a microscopic analysis of subtext inference. We operationally define subtext inference as the ability to uncover hidden information through deduction or by capturing pieces of information from the conversation. Subtext inference has two sub-abilities: information capture and information deduction. Information capture measures whether a spy notices and exploits critical hints inadvertently revealed by citizens. Since weak citizens generally leak such information unintentionally, information capture ability becomes essential for the spy to win the game against them. Since strong citizens rarely make such mistakes, it is highly challenging to observe information capture of a spy. On the other hand, information deduction mainly focuses on how effectively a spy infers a hidden location by combining various pieces of information. Since strong citizens generally conceal the key information, information deduction ability becomes essential for the spy to uncover information embodied in a citizen’s utterance.
@@ IV-B1 Metric design
To assess the spy’s ability to identify hidden intent, we propose four metrics based on their guessing behavior. These metrics range from relatively summative to more formative.
@@
  • Guess Success (GS) measures how much the guessing of a spy contributed to their victory. This metric focuses only on games where the spy attempted to guess, rather than all played games. We calculate GS by dividing the number of games in which the spy made a successful guess by the number of games in which guessing was attempted. A higher GS indicates that the spy has successfully identified hidden locations, demonstrating strong subtext inference skills.
  • Notice Rate (NR) indicates how well a spy noticed the hidden location across all games. It is calculated by computing the percentage of games in which the spy correctly detected the hidden location over the total number of games played. Unlike GS, which is measured only when the spy decides to reveal their identity publicly, NR can be computed across all games. This is because NR mainly considers whether the spy notices the hidden location, regardless of whether they decide to guess.
@@
  • Information Catching (IC) measures how often the spy successfully leverages exposed information to achieve a victory. Unlike NR, which only assesses the ability of the spy to identify the hidden location, IC evaluates the effectiveness in employing such information to win. We specifically focus on games where weak citizens directly exposed information about the hidden location instead of considering all the played games. Here, the exposure means a case where a citizen’s utterance literally contains the hidden location [^5] Qualitative analysis in Section V-B1 discusses the case.. We measured IC by computing the proportion of games in which the spy successfully guessed the hidden location from games where exposure occurred. Thus, a high IC score reflects the ability of the spy to achieve victory by effectively capturing and utilizing key information.
  • Information Deduction (ID) is similar to IC but it applies to games where exposure did not occur. This metric measures how well the spy combined pieces of information to determine the hidden location. To compute ID, we divided the number of games in which the spy successfully guessed the hidden location by the number of games without any exposure. Therefore, a high ID indicates that the spy successfully combined clues from the conversation and utilized them to achieve victory.
@@
We provide a multifaceted interpretation of subtext inference by jointly analyzing IC and ID metrics. Based on whether each metric is above or below its average, we classify model behavior into three combinations. The first combination, when both IC and ID are high, represents an ideal scenario in which the spy successfully identifies the hidden location regardless of the type of citizen. The second combination occurs when only one of the metrics, either IC or ID, is higher than the average. This suggests a relative lack of proficiency in the corresponding sub-ability. For instance, a model with high IC but low ID demonstrates strong information capture but weak information deduction. The final and least favorable combination for subtext inference is when both IC and ID are low. This combination indicates that the spy failed to infer the hidden location in most games.
Moreover, by combining GS and NR, we can explore whether the spy effectively incorporates the incoming information to maintain their guess. The ideal combination is when both GS and NR are high. This indicates that the spy consistently maintained a correct guess after identifying the hidden location, until the end of the game. In contrast, a high GS paired with a low NR suggests that while the spy struggled to predict the hidden location initially, they managed to maintain their guess once the correct location was identified. Meanwhile, the combination of low GS and high NR implies that although the spy successfully guessed the hidden location, the spy is prone to discarding the guess easily. The worst-case scenario is when both NR and GS are low, indicating a significant failure in both identifying and maintaining the correct hidden location.
@@ IV-B2 Result
Tables II (b) and III (b) present the measurements for subtext inference when playing against strong and weak citizens, respectively. The results indicate that GPT-4 has sufficient ability compared to the other models. GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and Llama (81.25, 50.0, 27.78, and 0%, respectively). The results in terms of NR show a similar trend: GPT-4 (61.90%), Gemini (33.33%), GPT-3.5 (28.57%), and Llama (9.52%). Finally, regarding ID, GPT-4 was the best model, followed by Gemini, GPT-3.5, and Llama (61.90, 30.0, 28.57, and 0%). In summary, GPT-4 demonstrated the best performance against strong citizens in subtext inference. Meanwhile, Llama showed the lowest performance against strong citizens.
Against weak citizens, GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and Llama (90.48, 85.71, 71.43, and 60%, respectively). NR displayed the same trend: GPT-4 (100%), Gemini (90.48%), GPT-3.5 (76.19%), and Llama (71.43%). Lastly, regarding IC, GPT-3.5 (92.86%) showed the best performance, differing in GS and NR, followed by GPT-4 (87.5%), Gemini (61.11%), and Llama (20.0%). Similar to the results against strong citizens, GPT-4 showcased good performance against weak citizens.
@@ IV-B3 Discussion
In this section, we derive insights using the proposed metrics by conducting a multifaceted analysis. This discussion encompasses three aspects: (1) a complementary relationship between IC and ID; (2) an exploration of the spy’s filtering behavior on incoming information; and (3) the need for formative metrics concerning subtext inference.
First, IC and ID are complementary metrics that enable multifaceted analysis of subtext inference. As aforementioned, these two metrics can be combined to form three distinct partitions, each representing different levels of subtext inference ability. Using these partitions, we can determine the best model possessing sufficient sub-abilities for subtext inference; In our analysis, GPT-4 demonstrated high scores in both IC and ID, with its ID being double that of the second-place model. Meanwhile, GPT-3.5 falls within the high-IC region but not the high-ID region. This indicates that while GPT-3.5 struggles to combine hints from the conversation to deduce the hidden location.
Second, by analyzing GS and NR, we can examine how the spy filters incoming information to maintain their guess. As suggested in the metric design, we can use four quadrants made by GS and NR. Among the four models, GPT-4 belongs to the first partition and achieves a high score on both metrics. This indicates that GPT-4 can filter useful information, reinforcing the belief of the spy in the hidden location. This could be a hint that GPT-4 has strong information deduction sub-ability. Meanwhile, belonging to different partitions can induce differences in such filtering behaviors. For example, although Gemini and GPT-3.5 exhibited similar NR scores against strong citizens, GS of Gemini was almost double that of GPT-3.5. This difference suggests that Gemini is more selective in using guess actions than GPT-3.5, as evidenced by the fact that Gemini used 14 guess actions regardless of citizens, whereas GPT-3.5 utilized 21 (against weak) and 18 (against strong) guess actions.
@@
Finally, our metrics regarding subtext inference enable more detailed analyses of the factors contributing to the scores of simple metrics. Let us consider the simple metric, Winning Rate (WR), on Gemini models. Against strong citizens, Gemini showed a similar WR to Llama. Meanwhile, against weak citizens, Gemini exhibited the same WR as GPT-4. However, these observations are somewhat shallow when we consider our metrics, as the three models display different subtext inference abilities. For example, against strong citizens, Gemini had a higher ID score than Llama in that Llama completely failed to combine clues about the hidden location. Similarly, against weak citizens, we can conclude that GPT-4 is slightly better than Gemini because of its higher GS and NR scores. Thus, proposed metrics can reveal finer differences between LLMs than existing metrics.
In addition, such a detailed analysis revealed that GPT-3.5 is overly confident in its deduction results. As illustrated in Tables II and III, the spy with GPT-3.5 demonstrated relatively lower NR scores; however, the spy used guessing actions more frequently. That is, the spy rushed into making guesses even when their inference about the hidden location was incorrect. One possible explanation is that GPT-3.5 may treat its own inference as hard evidence, as we will discuss in Section V. We suspect that this mechanism may have contributed to GPT-3.5 achieving the lowest WR score, despite its relatively good subtext inference ability.
IV-C Deceptive Control
@@
In addition to subtext inference, we designed four metrics for the multifaceted analysis of deceptive control. Here, we operationally define deceptive control as the ability to blend in with others by controlling information. Deceptive control encompasses two sub-abilities: evading suspicion and dispersing suspicion. Evading suspicion mainly considers how well the spy avoids being suspected by others. It is quantified by the number of suspicions directed at the spy. On the other hand, dispersing suspicion primarily considers how difficult the spy makes it for all citizens to converge on a single suspect. It is quantified by the probability distribution of suspicions. To estimate the number and distribution of suspicions, we utilized logs of Day/Final votes. The following subsections describe the design of the four metrics and their results. The results indicate that GPT-3.5 is the most powerful spy player, followed by GPT-4, Gemini, and Llama.
@@ IV-C1 Metric design
  • Caught Rate (CR) is how frequently the spy is caught by citizens in the Day/Final vote. We compute this by dividing the number of games in which the spy was caught by the number of games lost. Thus, a high CR indicates that the majority of citizens easily identified the spy player during the game. Note that this is a summative metric for the spy’s deceptive control ability because it only considers the final state of the game.
  • Vote Rate (VR) is the proportion of suspicions about the spy across all votes that occurred. We measure this by computing the percentage of votes the spy received from citizens across the total number of votes, excluding the spy’s votes. Therefore, a high VR indicates that the spy failed to evade suspicion during the game.
  • Vote Entropy (VE) is how scattered the suspicions about the spy are in the game. We measure this by computing the entropy of the votes that citizens casted for each game and averaging them. Note that in SpyGame, a player’s suspicion cannot affect another player’s suspicion because they cannot have a discussion together. Thus, when a spy successfully blends with citizens, citizens have difficulty narrowing down the number of spy candidates. Thus, the value of VE increases.
We devised a multifaceted interpretation of deceptive control by using VR and VE simultaneously. By partitioning VR and VE by their average values, we derived four combinations. First, the most balanced combination was a low VR and high VE. This is an ideal combination because the spy can successfully evade others’ suspicions and lead citizens to discourage each other from narrowing down spy candidates. Second, models in which VR and VE are both low or high lack one of the two sub-abilities. Finally, the worst combination occurs when VR is high and VE is low. This combination indicates that citizens’ suspicions were concentrated on the spy and that the spy failed to undermine suspicions.
@@ IV-C2 Result
Tables II (c) and III (c) elaborate measurements of deceptive control when playing against strong and weak citizens, respectively. Note that some missing values for VR and VE indicate that the spy guessed the hidden location for all games, and no votes were conducted.
The results indicate that the variants of GPT have sufficient ability compared to other models. Against strong citizens, GPT-3.5 was the best model in terms of CR, followed by Gemini, GPT-4, and Llama (7.1, 22.2, 25.0, and 100%, respectively). The results for VR showed a similar tendency: GPT-3.5 (25%), GPT-4 (33.3%), Gemini (35.7%), and Llama (45.2%). Finally, regarding VE, GPT-4 was the best model, followed by GPT-3.5, Llama, and Gemini (0.953, 0.730, 0.634, and 0.348, respectively). Interestingly, GPT-3.5 showed deceptive control performance similar to that of GPT-4, although GPT-4 showed much higher performance in other language-based tasks[1].
Against weak citizens, GPT variants showed good performance compared to others. Because GPT variants ended the game by guessing, citizens had no chance to find the spy. Thus, CR values for GPT-3.5 and GPT-4 were zero, and VR and VE could not be measured using these models. Other models showed a tendency similar to that against strong citizens. Regarding CR and VR, Gemini (0% and 7.1%, respectively) was better than Llama (60% and 13.5%). However, regarding VE, Gemini (0.420) performed worse than Llama (0.605), as shown by results against strong citizens.

### v3

@@ IV-B Analysis of Subtext Inference
Information capture measures whether a spy notices and exploits explicit clues inadvertently revealed by citizens. Since weak citizens generally leak such information unintentionally, the spy’s ability to capture information becomes essential to win the game against them. On the other hand, information deduction mainly focuses on how effectively a spy infers the hidden location by synthesizing implicit clues that citizens intentionally crafted to signal their identity to other citizens without revealing the location. Since strong citizens generally conceal key information, the skill of deducing information becomes essential for the spy to uncover the information embodied in a citizen’s utterance.
@@ IV-B1 Metric design
To evaluate how subtext inference contributes to the spy’s success and to diagnose the state of its two underlying skills, we propose three metrics.
TABLE IV: Multifaceted interpretation of subtext inference
Information Information Case Interpretation
Catching Deduction
Ideal High High The model effectively utilizes both sub-skills and demonstrates ideal subtext inference
Moderate High Low The model captures implicit clues well but struggles to synthesize them into a final deduction
Moderate Low High The model shows strong reasoning ability but fails to identify explicit clues from the conversation
Worst Low Low The model fails to capture and reason over subtext, indicating poor subtext inference ability
@@
  • Guess Success Rate (GSR) measures how much the guessing of a spy contributed to their victory. This metric focuses only on games where the spy attempted to guess, rather than all played games. We calculate GSR by dividing the number of games in which the spy made a successful guess by the number of games in which guessing was attempted. A higher GSR suggests that the spy is more capable of inferring subtextual information to reach out to hidden targets.
    $GSR=\frac{\#\text{ victory by }Guessing}{\#\text{ victory}}$ (3)
@@
  • Information Catching Rate (ICR) measures how effectively the spy exploits explicit clues directly stated by citizens to identify the hidden location. This metric focuses only on games in which exposure occurred rather than all games. We specifically focus on games where exposure occurred rather than all the played games. Here, the exposure refers to a case where a citizen’s utterance literally contains the hidden location [^5] Qualitative analysis in Section V-B1 discusses the case..
[…]
  • Information Deduction Rate (IDR) measures whether the spy can exploit implicit clues, which are scattered hints embedded across multiple citizen utterances, to infer the hidden location without exposure. This metric applies only to games where exposure did not occur, meaning the hidden location was never explicitly mentioned by any citizen. Unlike ICR, where success may depend on detecting an obvious clue, IDR reflects the spy’s capacity to synthesize indirect linguistic signals and reason under ambiguity. We compute IDR as the proportion of games without exposure in which the spy correctly guessed the location.
[…]
@@
As illustrated in Table IV, we provide a multifaceted interpretation of subtext inference by jointly analyzing ICR and IDR metrics. Based on whether each metric is above or below its average, we classify model behavior into three combinations. The first combination, when both ICR and IDR are high, represents an ideal scenario in which the spy successfully identifies the hidden location regardless of the type of citizen. The second combination occurs when only one of the metrics, either ICR or IDR, exceeds the average. This suggests a relative lack of proficiency in the corresponding skills. For instance, a model with high ICR but low IDR demonstrates strong information capture but weak information deduction. The final and least favorable combination for subtext inference is when both ICR and IDR are low. This combination indicates that the spy failed to infer the hidden location in most games. This categorization allows us to diagnose which skill each model lacks, providing insight into how the model could be improved to more effectively support obscured communication.
@@ IV-B2 Result
Tables V (b) and VI (b) present the measurements for subtext inference when playing against strong and weak citizens, respectively. When playing against strong citizens, we found there were statistically significant differences between LLMs in all metrics regarding the subtext inference, including Guess Success Rate (GSR, $\chi^{2}=8.68$, $p<0.05$) and Information Deduction Rate (IDR, $\chi^{2}=18.02$, $p<0.001$). Specifically, GPT-4 was the best model in terms of GSR, followed by Gemini, GPT-3.5, and Llama (81.25, 50.00, 31.57, and 0%). This result indicates that GPT-4 has sufficient ability compared to other models in achieving victory by utilizing collected information during the games. Similarly, GPT-4 also exhibited the best score in IDR, followed by Gemini, GPT-3.5, and Llama (61.90, 30.0, 28.57, and 0%). This suggests that GPT-4 possesses a strong ability to leverage implicit clues to reach out to the hidden location.
Against weak citizens, while the overall ranking trend among models for the Guess Success Rate (GSR) remained consistent, no statistically significant differences were observed. We attribute this to the fact that weak citizens frequently revealed explicit clues that allowed the spy to easily infer the hidden location. In contrast, we found statistically significant differences between the tested models in Information Catching Rate (ICR, $\chi^{2}=21.83$, $p<0.001$). Specifically, GPT-3.5 outperformed other models (92.86%), followed by GPT-4 (87.50%), Gemini (61.11%), and Llama (20.00%). This suggests that GPT-3.5 is particularly effective at exploiting explicit clues often leaked by weak citizens.
@@ IV-B3 Discussion
TABLE V: Quantitative results for playing against strong citizens
(a) Prior Metric (b) Subtext inference (c) Deceptive Control
Winning Living Guess Information Information Caught Vote Vote
Model Rate Round Success Rate Catching Rate Deduction Rate Rate Rate Entropy
Gemini 57.14 6.95 50.00 - 30.00 022.22 35.71 0.348
(14) (20) (9) (42)
GPT-3.5 33.33 3.71 31.57 - 28.57 007.14 25.00 0.730
(19) (21) (14) (12)
GPT-4 80.95 5.80 81.25 - 61.90 025.00 33.33 0.953
(16) (21) (4) (30)
Llama2 52.38 9.00 - - 00.00 100.00 46.03 0.657
(0) (19) (10) (126)
Total 55.95 6.36 53.06 - 30.86 037.83 40.95 0.642
Metrics whose lower values indicate better results.
Note that Information Catching cannot be   measured playing against strong citizens.
TABLE VI: Quantitative results for playing against weak citizens
(a) Prior Metric (b) Subtext inference (c) Deceptive control
Win Living Guess Information Information Caught Vote Vote
Model Rate Round Success Rate Catching Rate Deduction Rate Rate Rate Entropy
Gemini 90.47 5.28 85.71 61.11 - 00.00 07.14 0.420
(14) (18) (2) (42)
GPT-3.5 71.42 2.52 71.42 92.86 - 00.00 - -
(21) (14) (6) (0)
GPT-4 90.47 2.42 90.47 87.50 - 00.00 - -
(21) (16) (2) (0)
Llama2 76.19 7.85 60.00 20.00 - 60.00 15.62 0.605
(5) (15) (5) (96)
Total 82.14 4.52 80.32 65.07 - 20.00 13.04 0.548
Metrics whose lower values indicate better results.
Note that Information Deduction was cannot be   measured playing against weak citizens.
@@
In this section, we derive insights using the proposed metrics by conducting a multifaceted analysis. This discussion encompasses three aspects: (1) a complementary relationship between Information Catching Rate (ICR) and Information Deduction Rate (IDR); (2) revealing decision-making behaviors through comparative analysis of metric-level disparities; and (3) the need for formative metrics concerning subtext inference.
First, ICR and IDR are complementary metrics that enable multifaceted analysis of subtext inference. By using the categorizations illustrated in Table IV, we can determine the best model possessing sufficient skills for subtext inference; In our analysis, GPT-4 exhibits strong performance in both ICR and IDR, with its IDR score being twice as high as that of the second-best model. In contrast, GPT-3.5 shows high ICR but low IDR, suggesting that while it is effective at capturing explicit clues, it struggles to infer the hidden location when only subtle, implicit cues are available. This observation suggests that for models exhibiting a similar profile–high ICR but low IDR–future efforts to improve their ability to support obscured communication should focus on enhancing their capacity to infer hidden meanings from subtle contextual cues, rather than relying on explicit exposures.
Second, our metrics reveal a discrepancy between LLMs’ ability to collect information and their decision-making thresholds for acting on the information. While tested models differ significantly in their ability to gather information (ICR and IDR), their thresholds for acting appear more consistent. Specifically, although the maximum gaps in GSR across models are 49.68 and 30.47% against weak and strong citizens, respectively, the corresponding gaps in ICR and IDR are much larger: 72.86 and 61.90%, respectively. In addition to these numeric differences, statistical tests further confirm that model differences in ICR and IDR are more pronounced than in GSR. This suggests that although LLMs vary greatly in how well they capture or synthesize clues from context, they tend to behave similarly when deciding whether enough evidence has been gathered to initiate a guess. In other words, the threshold for taking action based on inferred information appears more uniformly developed than the underlying abilities to process that information.
@@
Finally, our metrics regarding subtext inference enable more detailed analyses of the factors contributing to the scores of simple metrics. Let us consider the simple metric, Winning Rate (WR), on Gemini models. Against strong citizens, Gemini showed a similar WR to Llama. Meanwhile, against weak citizens, Gemini exhibited the same numerical WR score as GPT-4. However, these observations are somewhat shallow when considering our metrics, as the three models display different subtext inference abilities. For example, against strong citizens, Gemini had a higher IDR score than Llama, as Llama completely failed to synthesize implicit clues about the hidden location. Similarly, against weak citizens, we can conclude that GPT-4 is better than Gemini because of its higher ICR score. Thus, proposed metrics can reveal finer differences between LLMs than existing metrics.
IV-C Analysis of Deceptive Control
@@
In addition to subtext inference, we designed three metrics for the second core aspect, deceptive control. This aspect concerns the model’s ability to blend in with others by controlling information. Deceptive control encompasses two skills: evading suspicion and dispersing suspicion. Evading suspicion mainly involves how well the spy avoids being suspected by others. It is quantified by the number of suspicions directed at the spy. On the other hand, dispersing suspicion primarily considers how difficult the spy makes it for all citizens to converge on a single suspect. It is quantified by the probability distribution of suspicions. To estimate the number and distribution of suspicions, we utilized logs of Day/Final votes. The following subsections describe the design of the four metrics and their results.
@@ IV-C1 Metric design
  • Caught Rate (CR) measures how much the voting of citizens contributed to the spy’s failure. We compute this by dividing the number of games in which the spy was caught by the number of games the spy lost. Thus, a high CR indicates that the majority of citizens easily identified the spy player during the game. Note that this is a summative metric for the spy’s deceptive control ability because it only considers the final state of the game.
    $CR=\frac{\#\text{ defeat by }Voting}{\#\text{ defeat}}$ (6)
  • Vote Rate (VR) measures how much suspicion was concentrated on the spy during a game. We calculate it as the proportion of total votes that citizens cast against the spy in each game. Therefore, a high VR indicates that the spy failed to evade suspicion during the game.
[…]
  • Vote Entropy (VE) is how scattered the suspicions about the spy are in the game. We measure this by computing the entropy of the votes that citizens cast for each game and averaging them. Note that in SpyGame, a player’s suspicion cannot affect another player’s suspicion because they cannot have a discussion together. Thus, when a spy successfully blends with citizens, citizens have difficulty narrowing down the number of spy candidates. Thus, the value of VE increases.
    $VE=\frac{\sum(Entropy(\text{votes cast by citizens))}}{\#\text{games with }Voting}$ (8)
As illustrated in Table VII, we developed a multifaceted interpretation of deceptive control by jointly analyzing Vote Rate (VR) and Vote Entropy (VE). By partitioning both metrics based on their average values, we identified four representative combinations. First, the most desirable combination is low VR and high VE, which reflects an ideal form of deceptive control: the spy avoids being directly targeted while simultaneously inducing citizens to spread their suspicions across each other. Second, combinations where both VR and VE are either low or high suggest an imbalance or partial deficiency in the skills involved in deceptive control. For example, high values in both metrics indicate that although suspicion is spread among citizens, the spy remains a frequent target, which reflects only partial success in deception. Finally, the worst-case scenario is high VR and low VE, where the spy becomes the focal point of suspicion and fails to redirect or confuse citizens.
@@ IV-C2 Result
TABLE VII: Multifaceted interpretation of deceptive control
Vote Vote Case Interpretation
Rate Entropy
Ideal Low High The model successfully avoids direct suspicion while distributing suspicion among citizens
Moderate High High Although the model encourages suspicion to be spread, it still attracts frequent votes
Moderate Low Low The model avoids being voted, but fails to diffuse suspicion among citizens
Worst High Low The model becomes the primary target of suspicion and fails to spread suspicion across other players
Tables V (c) and VI (c) elaborate measurements of deceptive control when playing against strong and weak citizens, respectively. Note that some missing values for Vote Rate (VR) and Vote Entropy (VE) indicate that the spy guessed the hidden location for all games, and no votes were conducted.
Against playing strong citizens, we found statistically significant differences between LLMs in Caught Rate (CR, $\chi^{2}=23.25$, $p<0.001$) while no differences in VR or VE. Specifically, GPT-3.5 achieved the lowest CR (7.14%), followed by Gemini, GPT-4, and Llama (22.22, 25.00, and 100.00%). Since CR measures how much the voting of citizens contributed to the spy’s failure, this result suggests that GPT-3.5 was rarely identified as the spy when it lost. Meanwhile, VR and VE did not show statistically significant differences across models, although the overall ranking trends were similar to those of CR. We interpret the small differences in VR and VE as a reflection of how models with high CR, such as Llama, were often caught after receiving a minimal majority of votes in every voting. In other words, although they lost, they were not overwhelmingly targeted against strong citizens.
Against weak citizens, no statistically significant differences were observed across models in any of the deceptive control metrics. We attribute the lack of difference in CR to the overall high winning rates, which resulted in very few lost games and thus limited opportunities for the spy to be caught via votes. Due to the early termination of most games, voting occurred infrequently. As a result, VR and VE could not be measured for GPT variants. These observations suggest that, when playing against weak citizens, the spy’s success depends less on the ability to deceive opponents and more on the ability to exploit explicit clues to directly guess the hidden location.